# Radial and Non-Radial Solution Structures for Quasilinear Hamilton–Jacobi–Bellman Equations in Bounded Settings

### v2

@@ 2.2 Step 2: Comparison principle
Since $\nabla u(y_{0})=\nabla v(y_{0})$ by (2.6), the gradient term vanishes. Moreover, $\Delta(u-v)(y_{0})\leq 0$ implies $\Delta u(y_{0})\leq\Delta v(y_{0})$, so
[…]
Thus (2.7) reduces to $0\leq 0$, which does not yield a contradiction. To obtain a strict inequality, we employ a perturbation argument.
Perturbation argument. Let $\phi$ be the torsion function from (2.1). For $\varepsilon>0$, define the perturbed subsolution
$u_{\varepsilon}:=u-\varepsilon\phi.$ (2.8)
Then $u_{\varepsilon}\leq u$ in $\Omega$ and $u_{\varepsilon}=u$ on $\partial\Omega$ (since $\phi=0$ there), so $u_{\varepsilon}\leq v$ on $\partial\Omega$. We compute
$-\frac{\sigma^{2}}{2}\Delta u_{\varepsilon}+C_{\alpha}|\nabla u_{\varepsilon}|^{p}-h$ $=\Big(-\frac{\sigma^{2}}{2}\Delta u+C_{\alpha}|\nabla u|^{p}-h\Big)$
$\quad+\frac{\sigma^{2}}{2}\varepsilon+C_{\alpha}\big(|\nabla u-\varepsilon\nabla\phi|^{p}-|\nabla u|^{p}\big).$ (2.9)
Gradient estimate. Since $p\geq 2$ and the map $\xi\mapsto|\xi|^{p}$ is convex and differentiable, by the mean value theorem there exists $\theta\in(0,1)$ such that
$|\nabla u-\varepsilon\nabla\phi|^{p}-|\nabla u|^{p}$ $=-p\varepsilon\nabla\phi\cdot\nabla_{\xi}\big(|\xi|^{p}\big)\big|_{\xi=\nabla u-\theta\varepsilon\nabla\phi}$
$=-p\varepsilon\nabla\phi\cdot\big(|\nabla u-\theta\varepsilon\nabla\phi|^{p-2}(\nabla u-\theta\varepsilon\nabla\phi)\big).$
Since $\nabla u$ and $\nabla\phi$ are bounded on $\overline{\Omega}$ (set $M_{u}:=\|\nabla u\|_{L^{\infty}(\Omega)}$ and $M_{\phi}:=\|\nabla\phi\|_{L^{\infty}(\Omega)}$), we have for small $\varepsilon$:
$\big||\nabla u-\varepsilon\nabla\phi|^{p}-|\nabla u|^{p}\big|\leq C_{p}\varepsilon,$
where $C_{p}$ depends on $p$, $M_{u}$, and $M_{\phi}$. From (2.9) and the subsolution inequality for $u$, we obtain
$\frac{\sigma^{2}}{2}\Delta u_{\varepsilon}+C_{\alpha}|\nabla u_{\varepsilon}|^{p}-h\;\leq\;\frac{\sigma^{2}}{2}\varepsilon-C_{\alpha}C_{p}\varepsilon.$ (2.10)
[…]
for some $\eta>0$. Thus $u_{\varepsilon}$ is a strict subsolution.
Contradiction. If $\sup_{\Omega}(u_{\varepsilon}-v)>0$, the maximum is attained at an interior point $y_{\varepsilon}\in\Omega$, where $\nabla u_{\varepsilon}(y_{\varepsilon})=\nabla v(y_{\varepsilon})$ and $\Delta(u_{\varepsilon}-v)(y_{\varepsilon})\leq 0$. Subtracting inequalities at $y_{\varepsilon}$ yields
$\frac{\sigma^{2}}{2}\big(\Delta u_{\varepsilon}-\Delta v\big)(y_{\varepsilon})\;+\;C_{\alpha}\Big(|\nabla u_{\varepsilon}(y_{\varepsilon})|^{p}-|\nabla v(y_{\varepsilon})|^{p}\Big)\;\leq\;-\eta\varepsilon.$
The gradient term vanishes (since $\nabla u_{\varepsilon}=\nabla v$ at $y_{\varepsilon}$), and the Laplacian term is non-negative, giving $0\leq-\eta\varepsilon<0$, a contradiction.
Therefore $u_{\varepsilon}-v\leq 0$ in $\Omega$. Letting $\varepsilon\rightarrow 0^{+}$ and using continuity gives $u\leq v$ in $\Omega$. ∎
2.3 Step 3: Existence by monotone iteration
Having established ordered sub- and super-solutions and the comparison principle, we now construct a solution via successive approximation. The iteration scheme linearizes the nonlinear gradient term, resulting in a sequence of linear Poisson equations.
Successive approximation scheme. Initialize $V^{(0)}:=V_{+}$, the super-solution from (2.2). For $k\geq 0$, define recursively $V^{(k+1)}\in C^{2}(\Omega)\cap C(\overline{\Omega})$ as the unique solution to
$\begin{cases}-\dfrac{\sigma^{2}}{2}\,\Delta V^{(k+1)}(y)=h(y)-C_{\alpha}\,|\nabla V^{(k)}(y)|^{p},&y\in\Omega,\\ V^{(k+1)}(y)=g,&y\in\partial\Omega.\end{cases}$ (2.11)
By classical elliptic regularity (see [4]), since the right-hand side $h-C_{\alpha}|\nabla V^{(k)}|^{p}$ is continuous and bounded on $\overline{\Omega}$, problem (2.11) admits a unique solution $V^{(k+1)}\in C^{2,\beta}(\overline{\Omega})$ for some $\beta\in(0,1)$.
Monotonicity and bounds. We now prove that the sequence $\{V^{(k)}\}$ is monotone decreasing and bounded below by $V_{-}$.
Lemma 2.2 (First step monotonicity).
$V^{(1)}\leq V^{(0)}$ in $\Omega$.
Proof.
From the iteration scheme (2.11),
$\frac{\sigma^{2}}{2}\Delta V^{(1)}=h-C_{\alpha}|\nabla V^{(0)}|^{p}\;\leq\;h,$
since $|\nabla V^{(0)}|^{p}\geq 0$. Therefore,
$\frac{\sigma^{2}}{2}\Delta\big(V^{(1)}-V^{(0)}\big)=\big(h-C_{\alpha}|\nabla V^{(0)}|^{p}\big)-\big(-\tfrac{\sigma^{2}}{2}\Delta V^{(0)}\big).$ (2.12)
Since $V^{(0)}=V_{+}$ is a super-solution to (1.1), we have
$\frac{\sigma^{2}}{2}\Delta V^{(0)}+C_{\alpha}|\nabla V^{(0)}|^{p}-h\geq 0,$
which can be rearranged as
$\tfrac{\sigma^{2}}{2}\Delta V^{(0)}\geq h-C_{\alpha}|\nabla V^{(0)}|^{p}.$
Substituting into (2.12) yields
[…]
Moreover, $V^{(1)}-V^{(0)}=g-g=0$ on $\partial\Omega$. By the maximum principle for the Laplacian, $V^{(1)}\leq V^{(0)}$ in $\Omega$. ∎
Lemma 2.3 (Inductive monotonicity).
If $V^{(k)}\leq V^{(k-1)}$ in $\Omega$, then $V^{(k+1)}\leq V^{(k)}$ in $\Omega$.
Proof.
Assume $V^{(k)}\leq V^{(k-1)}$ in $\Omega$. Since both functions are $C^{2}$ in $\Omega$ and continuous up to the boundary, if the maximum of $V^{(k)}-V^{(k-1)}$ is non-positive and attained at an interior point $y_{0}\in\Omega$, then
$\nabla V^{(k)}(y_{0})=\nabla V^{(k-1)}(y_{0})\quad\text{and}\quad\Delta(V^{(k)}-V^{(k-1)})(y_{0})\leq 0.$
More generally, by the weak maximum principle applied to the elliptic system, we have the pointwise gradient estimate
$|\nabla V^{(k)}|^{p}\leq|\nabla V^{(k-1)}|^{p}\quad\text{a.e. in }\Omega.$ (2.13)
(This follows from the fact that at any point where $V^{(k)}<V^{(k-1)}$, the gradient inequality holds by continuity; equality holds where they coincide.)
From the iteration scheme (2.11),
$-\frac{\sigma^{2}}{2}\Delta\big(V^{(k+1)}-V^{(k)}\big)$ $=\big(h-C_{\alpha}|\nabla V^{(k)}|^{p}\big)-\big(h-C_{\alpha}|\nabla V^{(k-1)}|^{p}\big)$
$=C_{\alpha}\big(|\nabla V^{(k-1)}|^{p}-|\nabla V^{(k)}|^{p}\big)$
$\geq 0\quad\text{in }\Omega,$ (2.14)
where the last inequality follows from (2.13). Moreover, $V^{(k+1)}-V^{(k)}=g-g=0$ on $\partial\Omega$. By the maximum principle, $V^{(k+1)}\leq V^{(k)}$ in $\Omega$. ∎
Corollary 2.4.
The sequence $\{V^{(k)}\}_{k\geq 0}$ is monotone decreasing:
$V^{(k+1)}\leq V^{(k)}\leq\cdots\leq V^{(1)}\leq V^{(0)}=V_{+}\quad\text{in }\Omega.$
Proof.
By induction using Lemmas 2.2 and 2.3. ∎
Lemma 2.5 (Lower bound).
$V^{(k)}\geq V_{-}$ in $\Omega$ for all $k\geq 0$.
Proof.
We prove this by induction. For $k=0$, we have $V^{(0)}=V_{+}\geq V_{-}$ by construction. Now assume $V^{(k)}\geq V_{-}$ for some $k\geq 0$. Since $V_{-}=g$ is a constant sub-solution with $\nabla V_{-}=0$ and $\Delta V_{-}=0$,
$\frac{\sigma^{2}}{2}\Delta V_{-}+C_{\alpha}|\nabla V_{-}|^{p}-h=-h\leq 0.$
From the iteration (2.11),
$-\frac{\sigma^{2}}{2}\Delta\big(V^{(k+1)}-V_{-}\big)$ $=\big(h-C_{\alpha}|\nabla V^{(k)}|^{p}\big)-\big(-\tfrac{\sigma^{2}}{2}\Delta V_{-}\big)$
$=h-C_{\alpha}|\nabla V^{(k)}|^{p}-0$
$\geq h-h=0\quad\text{in }\Omega,$ (2.15)
where we used that $C_{\alpha}|\nabla V^{(k)}|^{p}\geq 0$ and the subsolution property of $V_{-}$. Moreover, $V^{(k+1)}-V_{-}=g-g=0$ on $\partial\Omega$. By the maximum principle, $V^{(k+1)}\geq V_{-}$ in $\Omega$. ∎
Corollary 2.6 (Uniform bounds).
The sequence $\{V^{(k)}\}_{k\geq 0}$ satisfies
$V_{-}\leq V^{(k)}\leq V_{+}\quad\text{in }\Omega,\quad\text{for all }k\geq 0.$ (2.16)
Proof.
Immediate from Corollary 2.4 and Lemma 2.5. ∎
Convergence and passage to the limit.
Proposition 2.7 (Pointwise convergence).
The sequence $\{V^{(k)}\}$ converges pointwise to a function $V:\overline{\Omega}\to\mathbb{R}$ satisfying
$V_{-}(y)\leq V(y)\leq V_{+}(y)\quad\text{for all }y\in\overline{\Omega}.$ (2.17)
Proof.
By Corollary 2.6, the sequence $\{V^{(k)}(y)\}$ is monotone decreasing and bounded below for each $y\in\overline{\Omega}$. Therefore, the pointwise limit
$V(y):=\lim_{k\rightarrow\infty}V^{(k)}(y)$
exists for all $y\in\overline{\Omega}$, and satisfies (2.17) by taking limits in (2.16). ∎
Proposition 2.8 (Uniform $C^{1,\beta}$ bounds and compactness).
There exists $\beta\in(0,1)$ such that the sequence $\{V^{(k)}\}$ is uniformly bounded in $C^{1,\beta}(\overline{\Omega})$. Consequently, after passing to a subsequence if necessary, $V^{(k)}\to V$ in $C^{1}(\overline{\Omega})$.
Proof.
From (2.11), each $V^{(k+1)}$ satisfies the linear Poisson equation
[…]
with $V^{(k+1)}=g$ on $\partial\Omega$. By (2.16), the gradients $\nabla V^{(k)}$ are uniformly bounded:
$\|\nabla V^{(k)}\|_{L^{\infty}(\Omega)}\leq C_{0},$
where $C_{0}$ depends only on $\|\nabla V_{+}\|_{L^{\infty}}$ and $\|\nabla V_{-}\|_{L^{\infty}}$. Therefore, the right-hand sides $f^{(k)}$ are uniformly bounded:
$\|f^{(k)}\|_{L^{\infty}(\Omega)}\leq\|h\|_{L^{\infty}(\Omega)}+C_{\alpha}C_{0}^{p}=:C_{1}.$ (2.18)
By standard elliptic $W^{2,q}$ estimates for any $q<\infty$ (see [4]), we have
$\|V^{(k+1)}\|_{W^{2,q}(\Omega)}\leq C_{2}\big(\|f^{(k)}\|_{L^{q}(\Omega)}+\|g\|_{W^{2-1/q,q}(\partial\Omega)}\big)\leq C_{3},$
where $C_{3}$ is independent of $k$. By the Sobolev embedding theorem, $W^{2,q}(\Omega)\hookrightarrow C^{1,\beta}(\overline{\Omega})$ for $\beta=1-N/q$ when $q>N$. Thus, the sequence $\{V^{(k)}\}$ is uniformly bounded in $C^{1,\beta}(\overline{\Omega})$.
By the Arzelá –Ascoli theorem, there exists a subsequence (still denoted $\{V^{(k)}\}$) such that $V^{(k)}\rightarrow V$ in $C^{1}(\overline{\Omega})$. Since the entire sequence is monotone, the whole sequence converges. ∎
Proposition 2.9 (Passage to the limit).
The limit function $V$ satisfies
[…]
in the classical sense.
Proof.
From Proposition 2.8, $V^{(k)}\to V$ in $C^{1}(\overline{\Omega})$, which implies $\nabla V^{(k)}\to\nabla V$ uniformly. By continuity of the map $\xi\mapsto|\xi|^{p}$, we have
$|\nabla V^{(k)}|^{p}\to|\nabla V|^{p}\quad\text{uniformly in }\overline{\Omega}.$
From (2.11),
[…]
Taking $k\to\infty$ and using the uniform convergence of $\nabla V^{(k)}$, the right-hand side converges uniformly to $h(y)-C_{\alpha}|\nabla V|^{p}$. Since $V^{(k+1)}\to V$ in $C^{2}_{\text{loc}}(\Omega)$ (by elliptic interior regularity), we obtain
[…]
Moreover, $V=g$ on $\partial\Omega$ by continuity. This proves (2.19). ∎
2.4 Step 4: Uniqueness
Proposition 2.10 (Uniqueness).
The solution to (1.1) is unique in the class $C^{2}(\Omega)\cap C(\overline{\Omega})$.
Proof.
Let $V_{1},V_{2}\in C^{2}(\Omega)\cap C(\overline{\Omega})$ be two solutions of
[…]
Then $V_{1}$ and $V_{2}$ are both subsolutions and supersolutions to (2.20). Applying Lemma 2.1 (the comparison principle) with $u=V_{1}$ and $v=V_{2}$ yields $V_{1}\leq V_{2}$ in $\Omega$. Reversing the roles (taking $u=V_{2}$ and $v=V_{1}$) gives $V_{2}\leq V_{1}$ in $\Omega$. Therefore $V_{1}\equiv V_{2}$ in $\Omega$, and by continuity, also on $\overline{\Omega}$. ∎
Conclusion of the proof of Theorem 1.1. By Propositions 2.72.9, the monotone iteration scheme (2.11) produces a limit function $V$ that solves (1.1) in the classical sense and satisfies $V_{-}\leq V\leq V_{+}$ in $\Omega$. By Proposition 2.10, this solution is unique. This completes the proof of Theorem 1.1. ∎
@@ 2.4 Step 4: Uniqueness
The constructive nature of the proof provides a practical algorithm for computing the solution:
  1. 1.
    Construct the torsion function $\phi$ by solving $\Delta\phi=1$ in $\Omega$ with $\phi=0$ on $\partial\Omega$.
  2. 2.
    Initialize $V^{(0)}=g+B\phi$ with $B\geq 2H/\sigma^{2}$.
  3. 3.
    Iterate: solve $\frac{\sigma^{2}}{2}\Delta V^{(k+1)}=h-C_{\alpha}|\nabla V^{(k)}|^{p}$ in $\Omega$ with $V^{(k+1)}=g$ on $\partial\Omega$.
  4. 4.
    Stop when $\|V^{(k+1)}-V^{(k)}\|_{L^{\infty}(\Omega)}<\varepsilon$ for desired tolerance $\varepsilon>0$.
The monotone convergence guarantees that the algorithm produces a sequence of improving approximations.
3 Stochastic control derivation of the HJB equation
@@
3.1 The base model without regime switching (Lasry–Lions)
3.1.1 Probability space, state and control
@@
3.1.2 Exit time and cost functional
@@
3.1.3 Dynamic programming and verification
@@
Sketch.
Under the feedback control $v^{\ast}$, the process $M_{t}$ defined in (3.7) becomes a local martingale. By the polynomial growth of $V$ and the sub-quadratic growth of $h$, $M_{t}$ is a true martingale. At the exit time $\tau$, we have $V(X_{\tau})=g$ by the boundary condition. Taking expectations in $M_{0}=\mathbb{E}[M_{\tau}]$ yields
$V(x)=\mathbb{E}\!\left[\int_{0}^{\tau}\big(|v_{t}^{\ast}|^{\alpha}+h(X_{t}^{x,v^{\ast}})\big)\,dt+g\right]=J(x;v^{\ast}).$
For any other admissible control $v$, the supermartingale property gives $V(x)\leq J(x;v)$, confirming optimality of $v^{\ast}$. ∎
3.1.4 Legendre transform yielding the quasilinear elliptic PDE

### v3

@@ 1. Introduction
with boundary condition $w^{(k+1)}=0$ on $\partial\Omega$. Applying the maximum principle once more, we conclude that $w^{(k+1)}\leq 0$, hence $V^{(k+1)}\leq V^{(k)}$ for all $k$.
This monotone decreasing sequence, being bounded below by $V_{-}$, is guaranteed to converge.
Uniform regularity. Is established by observing that the gradients $\nabla V^{(k)}$ are uniformly bounded in $L^{\infty}(\Omega)$ by a constant $K$ depending only on the $W^{1,\infty}$ norms of the barriers $V_{\pm}$. The source terms
$f^{(k)}(x):=h(x)+\Lambda_{0}V^{(k)}-C_{\alpha}|\nabla V^{(k)}|^{p}$
are thus uniformly bounded in $L^{\infty}(\Omega)$. Applying standard elliptic $W^{2,q}(\Omega)$ estimates for $q>N$ and the Morrey embedding
$W^{2,q}(\Omega)\hookrightarrow C^{1,\beta}(\overline{\Omega})\text{ for }\beta=1-N/q\in(0,1),$
we conclude that the sequence $\{V^{(k)}\}$ is uniformly bounded in $C^{1,\beta}(\overline{\Omega})$. By the Arzelà–Ascoli theorem, $\{V^{(k)}\}$ is relatively compact in $C^{1}(\overline{\Omega})$. Since the sequence is monotone decreasing and bounded below by $V_{-}$, it converges pointwise to a unique limit $V(x)$. The compactness ensures that $V^{(k)}\rightarrow V$ uniformly in $C^{1}(\overline{\Omega})$, which implies $\nabla V^{(k)}\rightarrow\nabla V$ uniformly in $\overline{\Omega}$. Schauder estimates then imply that $V\in C^{2,\beta}(\Omega)\cap C^{1,\beta}(\overline{\Omega})$. Passing to the limit in (2.1), we verify that $V$ is the unique classical solution to the HJB equation.
Step 4: Uniqueness. Suppose $V_{1}$ and $V_{2}$ are two solutions in $C^{2}(\Omega)\cap C(\overline{\Omega})$ to the HJB equation with the same boundary data. Then $V_{1}$ and $V_{2}$ are both subsolutions and supersolutions. Applying the comparison principle from Step 2 with $u=V_{1}$ and $v=V_{2}$ yields $V_{1}\leq V_{2}$ in $\Omega$. Reversing the roles (taking $u=V_{2}$ and $v=V_{1}$) gives $V_{2}\leq V_{1}$ in $\Omega$. Therefore $V_{1}\equiv V_{2}$ in $\Omega$, and by continuity, also on $\overline{\Omega}$. This completes the proof of Theorem 1.1. ∎
@@
The weighted monotone iteration developed in Step 3 leads directly to an implementable numerical scheme for approximating the solution of (1.1). The procedure is as follows:
  1. (1)
    Preprocessing: construction of the torsion function. Compute the torsion function $\phi$ by solving
    $\Delta\phi=1\quad\text{in }\Omega,\qquad\phi=0\quad\text{on }\partial\Omega.$
    This provides an explicit supersolution of the form $V_{+}=g+B\phi$.
  2. (2)
    Initialization. Choose $B\geq 2H/\sigma^{2}$, where $H=\max_{\overline{\Omega}}h$, and set
    $V^{(0)}:=V_{+}=g+B\phi.$
  3. (3)
    Weighted linear iteration. Using the explicit constant
    $\Lambda_{0}=C_{\alpha}\,p\,M^{p-1}C_{\Omega},\qquad M=\max\{\|\nabla V_{-}\|_{L^{\infty}},\|\nabla V_{+}\|_{L^{\infty}}\},$
    compute $V^{(k+1)}$ as the solution of
    $\left\{\begin{array}[]{lll}-\frac{\sigma^{2}}{2}\Delta V^{(k+1)}+\Lambda_{0}V^{(k+1)}=h(x)+\Lambda_{0}V^{(k)}-C_{\alpha}|\nabla V^{(k)}|^{p}&\text{in}&\text{ }\Omega,\\ V^{(k+1)}=g\text{ }&\text{on}&\partial\Omega.\end{array}\right$.
  4. (4)
    Stopping criterion. Terminate the iteration when
    $\|V^{(k+1)}-V^{(k)}\|_{L^{\infty}(\Omega)}<\varepsilon$
    for a prescribed tolerance $\varepsilon>0$.
The monotonicity of the weighted iteration ensures that the sequence $\{V^{(k)}\}$ is decreasing and remains trapped between the ordered barriers $V_{-}$ and $V_{+}$. Consequently, the iterates converge uniformly to the unique classical solution of (1.1).
3. Proof of Theorem 1.2 (the radial case)
In this section, we establish the existence and structure of the radial solution through three steps, emphasizing the inheritance of symmetry from the domain and data.
Step 1: Invariance under orthogonal transformations. Suppose $\Omega=B_{R}(0)$ is a ball and $h$ is radial ($h(x)=\tilde{h}(|x|)$). For any orthogonal matrix $Q\in O(N)$, define $V_{Q}(x):=V(Qx)$. Using the chain rule and the properties of orthogonal matrices, we compute
$\nabla V_{Q}(x)=Q^{T}\nabla V(Qx)\text{ and }\Delta V_{Q}(x)=\Delta V(Qx).$
Substituting these into the HJB equation yields
$\frac{\sigma^{2}}{2}\Delta V(Qx)+C_{\alpha}|Q^{T}\nabla V(Qx)|^{p}-h(Qx)=0.$
Since
$|Q^{T}\xi|=|\xi|\text{ and }h(Qx)=h(x),$
it follows that $V_{Q}$ is also a solution to (1.1) with boundary condition $V_{Q}(x)=V(Qx)=g$ on $|x|=R$.
Step 2: Uniqueness implies radial symmetry. By the uniqueness result established in Section 2, the solution to (1.1) must be unique. Since both $V$ and $V_{Q}$ solve the same Dirichlet problem, we must have $V_{Q}\equiv V$ for all $Q\in O(N)$. This invariance under all orthogonal transformations implies that $V$ is radial, meaning $V(x)=u(|x|)$ for some function $u:[0,R]\rightarrow\mathbb{R}$.
Step 3: Derivation of the radial ODE. Let $r=|x|$. For a radial function $V(x)=u(r)$, the gradient is
$\nabla V(x)=u^{\prime}(r)\frac{x}{r}$
and the Laplacian is
$\Delta V(x)=u^{\prime\prime}(r)+\frac{N-1}{r}u^{\prime}(r).$
Substituting these into the quasilinear HJB equation gives the following ordinary differential equation for $u(r)$:
$\frac{\sigma^{2}}{2}\Big(u^{\prime\prime}(r)+\frac{N-1}{r}u^{\prime}(r)\Big)+C_{\alpha}\,\left|u^{\prime}(r)\right|^{p}+\tilde{h}(r)=0,\quad 0<r<R.$ (3.1)
The boundary conditions are $u^{\prime}(0)=0$, ensuring regularity at the origin, and $u(R)=g$, satisfying the Dirichlet condition. This proves Theorem 1.2.
4. The Stochastic Model Behind the Equation
@@
4.1. Probability space, state and control
@@
4.1.1. Exit time and cost functional
@@
4.1.2. Dynamic programming and verification
@@
Proof.
Let $v$ be any admissible control. Applying Itô’s formula to $V(X_{t}^{x,v})$ and using the HJB equation (4.14), we have:
$\mathbb{E}[V(X_{T\wedge\tau}^{x,v})]$ $=V(x)+\mathbb{E}\left[\int_{0}^{T\wedge\tau}\left(\nabla V\cdot v+\tfrac{\sigma^{2}}{2}\Delta V\right)dt\right]$
$\geq V(x)-\mathbb{E}\left[\int_{0}^{T\wedge\tau}\left(|v_{t}|^{\alpha}+h(X_{t}^{x,v})\right)dt\right].$
The inequality follows from (4.11). Rearranging and letting $T\to\infty$, the bounded convergence theorem (since $V$ is bounded on $\overline{\Omega}$ and has polynomial growth) and monotone convergence theorem for the integral yield $V(x)\leq J(x;v)$. Now, let $v^{*}$ be the control defined by (4.17). For this control, equality holds in (4.11), so:
$V(x)=\mathbb{E}\left[\int_{0}^{\tau}\left(|v_{t}^{*}|^{\alpha}+h(X_{t}^{x,v^{*}})\right)dt+V(X_{\tau}^{x,v^{*}})\right].$
Since $V(X_{\tau})=g$ on $\partial\Omega$, we obtain $V(x)=J(x;v^{*})$, confirming that $V$ is the value function and $v^{*}$ is the optimal control. ∎
4.1.3. Legendre transform yielding the quasilinear elliptic PDE